# Memory-Efficient Sequential Pattern Mining with Hybrid Tries

### v2

@@ 4.3 Computational Time
Figure 6 gives the computational time of all algorithms. The computational time of any algorithm that exceeds system memory is irrelevant and not reported. In the small data sets, computational time is almost identical for all algorithms, with the exception of TreeMiner. For example, $\mathcal{BT}$Miner and $\mathcal{HT}$Miner show a 99% improvement in computation time compared to TreeMiner in the Kosarak data set, emphasizing the benefits of a less memory intensive trie model.
In the medium-size Twitch data set, $\mathcal{BT}$miner and $\mathcal{HT}$Miner show an average speedup of 69% for $\mathcal{BT}$Miner and 72% for $\mathcal{HT}$Miner over PrefixSpan. The advantage is greater for lower support thresholds, with a maximum improvement of 94% for $\mathcal{HT}$Miner. Results on the Spotify data set show a slight advantage for PrefixSpan, especially for lower support thresholds. Here, trie models are disadvantaged by the overhead of constructing their trie, which leads to a higher computational time of at most 30 seconds. The computational time favours trie models more as the support threshold decreases and more time is spent in the mining algorithm. TreeMiner is the fastest trie model with an average 15% speedup over $\mathcal{HT}$Miner. This shows the benefit of the richer information stored at the trie nodes of TreeMiner, albeit, with a higher memory usage.
@@
The $\mathcal{BT}$Miner and $\mathcal{HT}$Miner algorithms show high computational time in large datasets. In the Criteo data set, both algorithms reach the imposed 36000 seconds time limits for support thresholds lower or equal to 6%. The computational time was considerably high for the Genome data set, where $\mathcal{HT}$Miner reached the time limit at all thresholds. In particular, we observed slower than usual mining speed in the Genome data set as memory consumption was near system limits, increasing the overhead of memory access by the CPU. This shows that it is time-efficient to reduce memory consumption even when operating within but close to system limits.
@@
Overall, we observe that $\mathcal{BT}$Miner and $\mathcal{HT}$Miner are also faster than the state of the art, especially for larger data sets. Note that this is generally not the case for memory-efficient algorithms, as memory and time efficiency are often a trade-off. We conclude that $\mathcal{HT}$Miner is superior to all other SPM alternatives, providing considerably higher memory efficiency, often coupled with higher time efficiency.
@@ 5 Conclusion
This paper develops a memory-efficient SPM algorithm using trie models of the data set. Our methodology involves a new binary trie model $\mathcal{BT}$ that stores minimal information at its nodes to compactly represent the data set in memory. We show that this compact representation is sufficient for the subsequent SPM task, and develop a novel algorithm $\mathcal{BT}$Miner to mine all sequential patterns. We build on our trie model to develop a hybrid model $\mathcal{HT}$ that models prefixes with high overlap using a trie model, and postfixes with low overlap using a vector model. We integrate $\mathcal{BT}$Miner and prefix projection to develop $\mathcal{HT}$Miner, which can effectively mine $\mathcal{HT}$. We proved that $\mathcal{HT}$Miner is always smaller and more memory efficient than pure vector or trie models and their corresponding algorithms.
Numerical results on real-life test instances showed that on small and medium data sets, $\mathcal{HT}$Miner provides, on average, 86% and 91% memory savings compared to the state-of-the-art vector and trie model of the literature, respectively. Furthermore, $\mathcal{HT}$Miner was shown to be the only SPM algorithm capable of mining large data sets within 256GB of system memory. While memory efficiency is often a trade-off with time efficiency, $\mathcal{HT}$Miner also showed lower computational time than the state of the art, with an average improvement of 23% over vector models and 59% over trie models.

### v3

@@ 4.3 Computational Time
Figure 6 gives the computational time of all algorithms. The computational time of any algorithm that exceeds system memory is irrelevant and not reported.
In the small data set Kosarak, TreeMiner is the slowest algorithm, especially for higher support thresholds. This is mainly due to the time required to populate its trie labels, which turned out to be costly in both time and memory efficiency. All other algorithms are mostly the same for higher support thresholds—where fewer patterns are mined. However, the improvements of $\mathcal{HT}$Miner and $\mathcal{BT}$Miner over PrefixSpan become larger for smaller thresholds, with more than a 60% speedup (approximately 800 seconds) for both algorithms. Both algorithms show an average improvement of 95% over TreeMiner, emphasizing the benefits in computational time provided by a less memory intensive trie model.
For the small data set MSNBC, all algorithms are mostly similar in time efficiency on higher support thresholds, but significantly different on lower support thresholds. This indicates that all algorithms are similar in constructing their data set representation, but not in mining it. PrefixSpan and $\mathcal{BT}$Miner require approximately 11,000 seconds to mine the MSNBC data set at the lowest support threshold (mining more than 100,000,000 patterns). In contrast, TreeMiner and $\mathcal{HT}$Miner show close to an order of magnitude faster mining, requiring approximately 1,800 seconds. The faster mining of TreeMiner shows the possible benefits of a richer set of trie labels, which increases mining efficiency when it can be efficiently modeled and fitted into memory. The faster mining of $\mathcal{HT}$Miner shows the benefit of a hybrid data structure in the mining process—equaling the benefits provided by richer trie labels in time efficiency while consuming less memory.
In the medium-size Twitch data set, $\mathcal{BT}$miner and $\mathcal{HT}$Miner show an average speedup of 50% over PrefixSpan. Both the PrefixSpan and $\mathcal{BT}$Miner reach the imposed time limit for the lowest support threshold, where PrefixSpan finds 15% of total patterns and $\mathcal{BT}$Miner finds 65% of total patterns. $\mathcal{HT}$Miner can mine the entire data set in approximately 9,000 seconds, showing at least a 75% speedup. These show the benefits of mining trie paths that model multiple sequences over mining vector representations which model each sequence separately. In particular, the mining algorithm can mine multiple sequences in a pass over a single trie path $({\bf{r}},\dots,n)$, which requires $f_{n}$ more passes in vector models to mine the same sequences. When $f_{n}$ is lower, such as in the postfixes of the longer sequences of the MSNBC data set, $\mathcal{HT}$Miner becomes more time-efficient than $\mathcal{BT}$Miner.
Results on the Spotify data set show similar computational time on all algorithms, with at most a $[20-40]$ seconds difference. TreeMiner is slightly faster (20 seconds) in higher support thresholds, while $\mathcal{HT}$Miner and $\mathcal{BT}$Miner are slightly faster (10 seconds) for lower support thresholds. PrefixSpan is consistently slower (5-40 seconds), with the difference highest for small support thresholds. This is mainly to the high compression of trie models given the low number of events $|\mathbb{E}|$ in this data set. Interestingly, $\mathcal{HT}$Miner and $\mathcal{BT}$Miner stay relatively constant in computational time over all support thresholds. This shows that most patterns are shorter in length, and demonstrates the efficiency of using ancestor labels in fast pruning of pattern extensions.
@@
The $\mathcal{BT}$Miner and $\mathcal{HT}$Miner algorithms show high computational time in large data sets. In the Criteo data set, both algorithms reach the imposed 36,000 seconds time limits for support thresholds lower or equal to 6%. The computational time was considerably high for the Genome data set, where $\mathcal{HT}$Miner reached the time limit at all thresholds. In particular, we observed slower than usual mining speed in the Genome data set as memory consumption was near system limits, increasing the overhead of memory access by the CPU. This shows that it is time-efficient to reduce memory consumption even when operating within—but close to—system limits.
@@
Overall, we observe that $\mathcal{HT}$Miner is also faster than the state of the art, especially for larger data sets. Note that this is generally not the case for memory-efficient algorithms, as memory and time efficiency are often a trade-off. For both time and memory efficiency, we thus conclude that $\mathcal{HT}$Miner is superior to all other SPM alternatives, providing considerably higher memory efficiency and often coupled with higher time efficiency.
@@ 5 Conclusion
This paper develops a memory-efficient SPM algorithm using trie models of the data set. Our methodology involves a new binary trie model $\mathcal{BT}$ that stores minimal information at its nodes to compactly represent the data set in memory. We show that this compact representation is sufficient for the subsequent SPM task, and develop a novel $\mathcal{BT}$Miner algorithm to mine all sequential patterns. We build on our trie model to develop a hybrid model $\mathcal{HT}$ that models prefixes with high overlap using a trie model, and postfixes with low overlap using a vector model. We integrate $\mathcal{BT}$Miner and prefix-projection to develop $\mathcal{HT}$Miner, which can effectively mine $\mathcal{HT}$. We proved that $\mathcal{HT}$Miner is always smaller and more memory efficient than pure vector or trie models and their corresponding algorithms.
Numerical results on real-life test instances showed that on small and medium data sets, $\mathcal{HT}$Miner provides, on average, 79% (approximately 5GB) and 90% (approximately 19GB) memory savings compared to the state-of-the-art vector and trie models, respectively. Furthermore, $\mathcal{HT}$Miner was shown to be the only SPM algorithm capable of mining large data sets within 256GB of system memory, potentially saving 1.7TB in memory consumption. While memory efficiency is often a trade-off with time efficiency, $\mathcal{HT}$Miner also showed lower computational time than the state of the art, with an average improvement of 43% over vector models and 54% over trie models.